OpenStudy (grace111):

(1) - Karl has £5. He spends 1/4 of his money . How much does he spend ? (2) - Asif is saving for a toy that costs £4 . He saves 40p a week . For how many weeks does he need to spend ? (3) - An MP3 download costs 80p . How many downloads can you get for £2? Thanks ... keep in mind will reward ! ^.^

OpenStudy (anonymous):

(1) money spent \[\frac{ 5 }{ 4 }=1.25\] thus karl has spend 1.25 pounds and karl saves 5-1.25=3.75 pounds (2)4 pounds=400p Time required to save 40p=1 week =>Time required to save 1p= 1/40 week =>Time required to save 400p(or 4 pounds)=(1/40)*400=10 weeks (3) 80p can give you amount of downloads - 1 1p can give you amount of downloads - 1/80 therefore 200p(or 2 pounds) can give you amount of downlodas - (1/80)*200 =20/8=5/2=2.5 downloads However it doesn't make sense how you can download in a fractional amount such as 2.5 Thus you can get 2 downloads for 160p with 40p remaining with you(you cannot download half an mp3 for 40p!) NOTE:It is important for your money to be either in completely pounds or completely pennies or your question will go wrong, notice how i've converted 2pounds into 200p and 4 pounds into 400p

OpenStudy (grace111):

thx so much :)